Two particles start moving from the same point along the same straight line. The first moves with constant velocity v and the second moves with acceleration :a'. During the time that elapses before the second catches the first, the greatest distance between the particles is...  Pls
Explain each step

Grade:11

1 Answers

Khimraj
5 years ago
let the time at which second catches first is t
then
vt = ½*a*t2
then
t = 2v/a
distance between them x = vt – ½*a*t2
for maximum x
dx/dt = 0
v -at = 0
t = v/a
then x = v2/a  – v2/2a = v2/2a
Hope it clears.